OpenStudy (anonymous):

The table below shows the relationship between years and average home price. Year 1990 1991 1992 1993 1994 1995 1996 Average Home Price 120,000 122,000 122,500 123,000 125,000 128,000 128,500 What is the equation of the line of best fit? (let x = the number of years after 1990)

OpenStudy (anonymous):

for the line of best fit, i would take the first point and the last point, which is year 1990, 1996 so you got two points (0, 120000) and (6, 128500) slope = (128500-120000)/6 = 1416.7 substitute a point in y = 1416.7x + b 120000 = b. y = 1416.7x + 120000

OpenStudy (anonymous):

Using the line of best fit, what would be the average price of a house in 2015?

OpenStudy (anonymous):

2015 - 1990 = 25 so substitute x = 25 y=155417.5